let z1=x+iy and z2=a+ib.suppose z1=z2. Which one of the following is/are true? Explain
a)x=a and y=b
b)the arguments of z1 and z2 differ by a multiple of 2pi
1
Expert's answer
2020-09-28T19:32:16-0400

a)

Definition of equality of two complex numbers:

Let "z_1 = x +iy, z_2 = a + ib"

"z_1 = z_2 \leftrightarrow Rez_1 = Rez_2, Imz_1 = Imz_2"


In our case:

"Rez_1 = x, Rez_2 = a \implies x =a"

"Imz_1 = y, Imz_2 = b \implies y = b"

So this following is true.


b)

The argument of complex number is angle between X axis and vector representing the complex number. Also each nonzero complex number has an infinite number of arguments that differ by a multiple of "2\pi k", where "k" - is any integer. Since z1 = z2, therefore this is the same number, it means their arguments differ by 2pi and second following is true.